I need to translate this word problem into an equation:
Price of a calculator. Evelyn paid $89.25, including 5% tax, for her graphing calculator. How much did the calculator itself cost?

Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

Price of a calculator. Evelyn paid $89.25, including 5% tax, for her graphing calculator. How much did the calculator itself cost?
--------------
Let "x" be the cost of the calculator.
EQUATION:
Cost = x + 0.05x
--------
89.25 = (1.05)x
x = $85.00
=================
Cheers,
Stan H.

Answer by edjones(8007)

x=original cost
x+.05x=89.25
1.05x=89.25
divide 1.05 into both sides. x=89.25/1.05=$85.00
Ed